### v1

# Advanced Vision Transformers and Open-Set Learning for Robust Mosquito Classification: A Novel Approach to Entomological Studies

Abstract
Mosquito-related diseases pose a significant threat to global public health, necessitating efficient and accurate mosquito classification for effective surveillance and control. This work presents an innovative approach to mosquito classification by leveraging state-of-the-art vision transformers and open-set learning techniques. A novel framework has been introduced that integrates Transformer-based deep learning models with comprehensive data augmentation and preprocessing methods, enabling robust and precise identification of ten mosquito species. The Swin Transformer model achieves the best performance for traditional closed-set learning with 99.80% accuracy and 0.998 F1 score. The lightweight MobileViT technique attains an almost similar accuracy of 98.90% with significantly reduced parameters and model complexities. Next, the applied deep learning models’ adaptability and generalizability in a static environment have been enhanced by using new classes of data samples during the inference stage that have not been included in the training set. The proposed framework’s ability to handle unseen classes like insects similar to mosquitoes, even humans, through open-set learning further enhances its practical applicability employing the OpenMax technique and Weibull distribution. The traditional CNN model, Xception, outperforms the latest transformer with higher accuracy and F1 score for open-set learning. The study’s findings highlight the transformative potential of advanced deep-learning architectures in entomology, providing a strong groundwork for future research and development in mosquito surveillance and vector control. The implications of this work extend beyond mosquito classification, offering valuable insights for broader ecological and environmental monitoring applications.
1 Introduction
Vector-borne diseases are illnesses resulting from pathogens, including bacteria, viruses, or parasites, which are transmitted to humans through the bites of vectors, especially mosquitoes. Mosquitoes are tiny insects well-known for their capacity to spread illness to people and animals through bites [1]. Mosquitoes are responsible for transmitting over 70 of vector-borne diseases globally [2]. There exist various debilitating diseases due to exposure to mosquito bites. For instance, Malaria is the deadliest disease transmitted by mosquitoes, accounting for more than 0.60 million deaths and 250 million cases worldwide in 2022. Dengue ranks second, with over 0.40 million deaths (mostly in Asia, particularly severe in Bangladesh) [3]. It occurs due to the bite of Aedes alobopictus and Aedes aegypti species of mosquitoes [4].
Recent advances in artificial intelligence (AI) and effective computer vision techniques provide significant progress in mosquito surveillance and monitoring, which is essential in the battle against vector-borne illnesses. Conventional identification and classification of mosquitoes include manual counting, entomological and microscopic observations, PCR and DNA tests after trapping from insecticide-treated nets, bednets, etc. These traditional strategies are inefficient, time-consuming and expensive due to the involvement of complex thermal cycling machines, medical entomologists and public health experts. Due to diminutive bodies, rapid movement, complex body landmarks, and striking similarity in the appearance of mosquitoes with that of other flies, the detection and classification of mosquitoes is exceptionally challenging. AI with visual sensors and computing facilities can simplify the difficult process of identifying mosquito species from trapped samples. For instance, deep learning models such as convolutional neural networks have illustrated impressive precision in categorizing mosquitoes captured via traps or cameras [5]. Additionally, AI-driven systems can analyze various data streams, including weather patterns and historical disease occurrences, to predict mosquito population dynamics and the risks of disease transmission. It is essential to accurately estimate mosquito vector population density on a large scale and insignificant processing time for preventing infectious diseases and deploying targeted interventions.
In this work, an AI-based automatic mosquito classification system has been developed using a comprehensive dataset of ten species, advanced vision Transformer models and open-set learning techniques. The primary contributions of this work are as follows:
A significant contribution of this research is to develop a comprehensive mosquito dataset by merging four publicly available datasets. This consolidated dataset encompasses ten distinct classes, including Aedes aegypti, Aedes albopictus, Anopheles albimanus, Anopheles arabiensis, Anopheles atroparvus, Anopheles coluzzi, Anopheles farauti, Anopheles freeborni, Anopheles stephensi, and Culex quinquefasciatus with 1749 sample images, including partially damaged and smashed images of mosquitoes.
Advanced data preprocessing and augmentation techniques are performed to balance the dataset.
Transfer learning and Transformer-based deep learning models, e.g., Xception, ViT, CvT, MobileViT and Swin, have been applied to classify the mosquito species into ten categories in traditional closed-set static environments.
Additionally, open-set learning techniques based on the OpenMax layer approach and Weibull distribution are integrated into the applied models, enabling them to identify and categorize images outside the predefined mosquito classes as “unknown.” The core principle involves calculating class-specific confidence scores based on the distance between a sample’s activation vector and the class mean activation vector, incorporating a Weibull distribution for robust uncertainty modeling.
The novelty of this work is to apply Transformer-based deep learning models and open-set learning techniques to classify mosquitoes using a comprehensive mosquito dataset of ten individual species.
2 Related work
The morphological characterization is important for accurate mosquito species classification. In recent times, significant work has been performed on AI-based automatic mosquito detection to aid taxonomic experts and public health professionals. For this reason, many researchers have shown their interest in classifying mosquitoes using deep learning approaches. Most of these used various pretrained classification and transformer models to classify mosquitoes through images and wingbeat sounds. Some recent articles on AI-based automatic mosquito detection have been briefly discussed in the following paragraphs.
Rustam et al. [6] used machine learning algorithms and innovative RIFS feature selection approaches to classify vector mosquito images for disease epidemiology. The authors classified two disease-spreading classes of mosquitoes, Aedes and Culex, using an open-source dataset from the IEEE data port. VGG16 model with the RIFS feature selection approach accomplished 99.2 accuracy with 0.986 F1 coefficient. A mosquito categorization scheme using transfer learning was developed by Isawasan et al. [7]. The authors classified two species of Aedes mosquito with a public database of 2400 images. They used five CNN models with transfer learning approaches (ResNet50, Xception, InceptionV3, MobileNetV2, and VGG16), and all the applied models achieved accuracies of more than 90 with VGG16 attaining a maximum accuracy of 98.25.
Siddiqui and Kayte [8] applied various deep learning frameworks for mosquito classification. They collected a dataset of six species of mosquitoes from website search engines employing web-scrapping techniques. The applied custom CNN and VGG16 models attained the maximum accuracies of 97.17 and 85.75, respectively. A comparative analysis of conventional and deep learning techniques for mosquito species categorization was investigated by Okayasu et al. [9]. The authors employed 14400 images from a public repository and three types of mosquito classes. The conventional SIFT method achieved the highest accuracy of 0.824. In contrast, the ResNet deep learning method had the highest of 0.955, demonstrating the effectiveness of deep classification in the categorization of mosquito species.
Lee et al. [10] investigated deep learning-integrated picture categorization for principal mosquito species in Korea. They used a dataset of 12 classes and eleven mosquito species comprising 7096 images. The Swin transformer hybrid model offered the best performance with an F1 score of 0.971 and 96.5 precision. Zhao et al. [11] initiated mosquito species identification utilizing various deep learning techniques. The dataset consists of seven genera and 17 species, with a total of 9900 mosquito images. The Swin transformer model accomplished the maximum accuracy, which is . YOLOv5 and DETR gave good competition to Swin transformers with an accuracy of and , respectively.
Omucheni et al. [12] used the Raman spectroscopy method and deep learning classification approaches to identify mosquito species. The authors used 397 mosquito images, including Anopheles gambiae and Anopheles arabiensis. The SVM model with a quadratic kernel performed best with 93 accuracy and 0.90 sensitivity. Rony et al. [13] investigated the identification of mosquito species by combining CNN and SVM with the help of a wingbeat frequency audio dataset. The applied models were trained using a large variety of data because of the acoustic sensitivity and the reliance of mosquito wing flapping on the weather and surroundings. Finally, the combined SVM and CNN model acquired 99.46 accuracy and 0.99 F1 score.
Yin et al. [14] used a deep learning-based pipeline for classifying and detecting mosquitoes based on their wingbeat noises. The data was collected from Medical Entomology of the Tropical Medicine facility of a Thailand university, which contained the wingbeat sound of four species of mosquitoes. The combined CNN and LSTM models secured the best performance with 61 precision and a 0.62 F1 score. After applying multiple audio preprocessing techniques, the classification accuracy improved to 90. An ensemble of various convolutional neural networks for new species recognition in mosquito species identification was done by Goodwin et al. [15]. The authors used a dataset containing 16 classes. Among the models they used, the CNN model Xception attained the best result for both closed-set and open-set. For the closed-set, it gave an accuracy and F1 score of 0.974 and 0.966, respectively. The Xception model’s accuracy and F1 score were reduced to 0.8906 and 0.7974 for the open-set novel class, respectively.
Pise and Patil [16] used a deep transfer learning framework for the multi-class classification of vector mosquito species of three classes. The dataset comprised 2,648 images, among which Aedes aegypti had 900 images, Anopheles stephensi had 540 images, and Culex Guinquefasciatus had 1208 images. GoogLeNet demonstrated superior performance, achieving a classification accuracy of 92.5% with feature extraction and 96% with fine-tuning of the corresponding hyperparameters.
The above studies demonstrate significant advancements in mosquito classification using various machine learning and deep learning approaches, with a particular emphasis on static environments. Several researchers have utilized traditional machine learning algorithms, convolutional neural networks (CNNs), and transformer models to classify mosquito species based on images and wingbeat sounds. In this constrained setup, training and testing of the models are performed with the same labels, leading to poor performance when the models encounter a new test class not present in the training samples. Due to mosquitoes’ close morphological characterization similarity with other flies and insects, such as double wings, antennae, tri-paired legs, and compound-structured eyes, open-set learning-based classification should be investigated. In this flexible, non-stationary situation, unknown testing samples may be present, which were not declared during the training stages. This realistic setup is very suitable for deployment in real applications, enhancing robustness and usability. This work explores the potential of vision transformer models with open-set learning for improving the accuracy and adaptability of mosquito classification systems, contributing to better disease control and epidemiology efforts.
3 Methodology
The proposed vision transformers and open-set learning for the mosquito classification system are described in detail in this section.
3.1 Dataset
3.1.1 Formation of the Dataset
The comprehensive dataset of this work has been constructed by combining various open-source image datasets utilized in related mosquito classification articles, e.g., image dataset of Aedes and Culex mosquito species [17], mosquito-on-human-skin [18], dataset of vector mosquito images [19] and malaria vector mosquito images [20]. The proposed dataset contains images of ten species of mosquitoes: two species of Aedes, seven of Anopheles and one Culex type. The dataset contains a total of 219 images of Aedes aegypti, 500 images of Aedes albopictus, 64 images of Anopheles albimanus, 193 images of Anopheles arabiensis, 36 images of Anopheles atroparvus, 74 images of Anopheles coluzzi, 57 images of Anopheles farauti, 97 images of Anopheles freeborni, nine images of Anopheles stephensi, and 500 images of Culex quinquefasciatus. The images of the two classes, Aedes albopictus and Culex quinquefasciatus, contain both whole and smashed/crushed mosquitoes, which can be described as partial mosquitoes, resulting in only parts of the insect being visible or intact. Table 1 represents the exact numbers of images for various mosquito classes. A sample image of each class is shown in Figure 1. Aedes is represented as Ae, Anopheles as An, and Culex as Cx in the caption of the figure.
| Type | Total |
| Aedes aegypti | 219 |
| Aedes albopictus | 500 |
| Anopheles albimanus | 64 |
| Anopheles arabiensis | 193 |
| Anopheles atroparvus | 36 |
| Anopheles coluzzi | 74 |
| Anopheles farauti | 57 |
| Anopheles freeborni | 97 |
| Anopheles stephensi | 9 |
| Culex quinquefasciatus | 500 |
| Total | 1749 |
3.1.2 Dataset Preprocessing
The data preprocessing approach is implemented to enhance the number of images by adding slightly altered imprints of previously existing data or creating new artificial information from the current data. It functions as a regularizer and helps prevent overfitting when a model is being trained [21]. Oversampling in data analysis is strongly connected to data augmentation techniques. In this work, additional images have been created from the original picture using standard data augmentation techniques, including rotation (20 degrees), width and height shifts (0.20), shear (0.20), zoom (0.20), horizontal flips, and nearest fill mode using ImageDataGenerator framwrok [22]. Figure 2 displays the number of augmented images of each mosquito class.
3.2 Model Training
This subsection describes the applied vision transformer-based deep learning models with their model architectures, open-set learning implementation and corresponding performance metrics.
3.2.1 ResNet50
ResNet50 is a deep convolutional neural network with 50 layers, utilizing residual learning to reduce the vanishing gradient problem through shortcut connections [23]. Training and validation datasets were loaded using flow_from_directory and image_dataset_from_directory with image dimensions 224224 and batch size 32. Excluding the top layer, the ResNet50 model was initialized with pre-trained weights from ImageNet. A new top layer was added consisting of a flattened layer, a dense network with 512 units and ReLU activation, a dropout layer (0.50), and a final dense layer with softmax activation. The model was compiled using the Adam optimizer (learning rate 0.0001), sparse categorical cross-entropy loss and 20 epochs. It is interesting to mention that, the Ray Tune Python library has been utilized in this work to select the optimized hyperparameters discussed above.
3.2.2 Xception
Xception enhances the Inception deep learning architecture by substituting the standard Inception modules with depthwise separable convolutions, leading to improved efficiency and performance [24]. The Xception model was initialized with pre-trained weights from ImageNet, excluding the top layer. Various hyperparameters involving the top layer for Xception models are found to be almost similar to ResNet50.
3.2.3 Vision Transformer (vit-base-patch16-224)
The Vision Transformer (ViT) applies the Transformer architecture to image analysis by treating image patches as sequences. This method enables the model to capture long-range dependencies within the image effectively. The specific variant utilized in this study is vit_base_patch16_224, which processes images in patches of 1616 [25]. Each image was resized to 256256 pixels for image preprocessing, center-cropped to 224224 pixels, randomly horizontally flipped, and normalized. The ViT model was instantiated using the timm library with pre-trained weights and was further adapted to classify different mosquito species, aligning with the research objectives. The model was trained using the Adam optimizer with a learning rate of 0.0001 and cross-entropy loss.
3.2.4 Swin Transformer (Swin-B and Swin-S)
Swin Transformers achieve efficiency by focusing their self-attention mechanism on localized image regions that shift across processing layers [26]. This design allows the model to efficiently handle long-range dependencies with a manageable computational complexity [27]. The Swin Transformer models are transferred to the computational device to ensure efficient processing. These setups leverage the pre-trained models to accelerate training and enhance the model’s ability to generalize from the dataset used for classifying mosquito species. Both models were trained using the Adam optimizer with a learning rate of 0.0001 and cross-entropy loss. The Swin-S model was trained for five epochs, while the Swin-B model was trained for ten epochs, with a batch size of 32. The architecture of the Swin Transformer used in this work is illustrated in Figure 3.
3.2.5 Convolutional Vision Transformer (CVT-13)
The Convolutional Vision Transformer integrates convolutional layers with vision transformer components, resulting in a model that efficiently leverages local and global features for image classification [28]. The training images underwent additional augmentation techniques, such as random horizontal flips, to introduce variability and improve the model’s robustness to different orientations and lighting conditions. In contrast, the test images were only resized and center-cropped to maintain the integrity of the test conditions. The CVT-13 model was initialized with pre-trained weights from the Microsoft model hub and tailored for the classification of mosquito species. A feature extractor was configured with AutoFeatureExtractor.from_pretrained(’microsoft/cvt-13’, do_rescale=False), allowing direct use of the model’s capabilities without resizing the images to the size initially used during the training of the model. The model, CvtForImageClassification.from_pretrained(’microsoft/cvt-13’, num_labels=len(train_dataset.classes), ignore_mismatched_sizes=True), was then transferred to the computational device. This setup ensures compatibility with the dataset’s specific class sizes and feature requirements while allowing for efficient computation. The architecture of the CVT-13 model employed in this study is depicted in Figure 4.
3.2.6 MobileViT
MobileViT integrates the efficiency of MobileNets approaches with the representational power of Vision Transformers, resulting in a lightweight yet robust model suitable for mobile and resource-constrained environments [29]. Data was resized to 256256, center-cropped to 224224, randomly horizontally flipped, and normalized. The MobileViT model was created using the transformers library with pre-trained weights, modified for mosquito species classification. The model was trained using the Adam optimizer (learning rate 0.0001) and cross-entropy loss for ten epochs with batch size 32. The architecture of the MobileViT model applied in this work is shown in Figure 5.
3.3 Open-set Learning Implementation
The dataset for open-set learning has been prepared through the following steps:
OpenMax Dataset preparation: Construct a test set of 100 unknown images (unseen during the training step) – bee (33), butterfly (34), and fly (33).
Use a trained model to calculate the Mean Activation Vector (MAV) and fit the Weibull distribution.
Create a pickle file.
Use the pickle file to set an OpenMax threshold.
3.3.1 Logits Extraction:
where and is the number of classes ( = 10).
3.3.2 Mean Activation Vector (MAV):
where are the logits of images that belong to class and is the number of such images.
3.3.3 Euclidean Distances:
The Euclidean distance [33] between each logit vector and its associated class MAV is computed as follows:
3.3.4 Weibull Distribution Fitting:
A Weibull distribution to the distances for each class is approximated as:
where , and are the shape, scale and location parameters, respectively.
3.3.5 Recalibrated Scores:
For each logit vector, the recalibrated scores [34] are obtained using the Weibull CDF as:
3.3.6 OpenMax Threshold:
If the maximum recalibrated score, defined in 5, is below a definite threshold, the inference sample is classified as “Unknown.” If not, the class is assigned with the highest recalibrated score [35, 36]:
The proposed open-set learning algorithm based on the OpenMax probability estimation applied in this work has been summarized in Algorithm 1.
The OpenMax algorithm processes input by accepting an activation vector , mean vectors , and Weibull parameters , along with , the number of leading classes to be adjusted. The algorithm starts by sorting to identify the classes with the highest activations and initializes scaling factors to 1 for all classes. For each of the top classes, it modifies the scaling factors based on the Euclidean distance between the class activation and its mean vector, adjusted by the Weibull shape and scale parameters. This scaling reduces the influence of activations that are far from the model’s training data, representing increased uncertainty. The modified activations are then computed by applying these scaling factors, and an additional activation is calculated for the unknown class, representing all reduced probabilities. Softmax is applied to these activations to compute the final class probabilities . The algorithm decides to reject the input if the predicted class is the unknown class or if its probability is lower than a set threshold , enhancing the model’s ability to deal with unknown or ambiguous inputs effectively.
3.4 Workflow
Figure 6 illustrates a workflow diagram for the proposed deep learning-based mosquito classification system involving data collection, preprocessing, model training, and evaluation. The process begins with data collection and merging various mosquito images, followed by data splitting and augmentation. It includes a decision point to check if the data is balanced, and if not, the data is balanced before compiling and training various models (ResNet50, Xception, ViT, Swin-S, Swin-B, CVT, and MobileViT). The training is followed by a check for the desired accuracy, with options to adjust the corresponding hyperparameters and re-evaluate the models. Once the desired accuracy is achieved, open-set learning is implemented employing the OpenMax dataset and Weibull distribution, and reports are generated.
4 Results and Discussion
This section represents the findings of this study for both closed and open-set learning for static and non-stationary environments, respectively. The first subsection contains mosquito image classification for closed-set. Next, the open-set learning results are illustrated for the applied deep learning techniques employing the OpenMax dataset and Weibull distribution. Finally, the proposed mosquito classification system has been compared with the related articles.
4.1 Results for Closed-set Learning
Table 2 presents the performance metrics for various models applied to closed-set learning in the proposed mosquito classification. The Swin-S Transformer technique achieves the highest overall accuracy at 99.80%, with corresponding high precision, recall, and F1 scores, indicating its superior performance. MobileViT achieves satisfactory performance with a high accuracy of 98.90% and an F1 score of 0.989, while utilizing significantly fewer parameters (5.58 million), making it an efficient model for mosquito classification. Conversely, ResNet50 shows the lowest accuracy at 94.50%, but still demonstrates reliable performance with a precision of 0.937 and a recall of 0.933. These results highlight the effectiveness of transformer-based models, particularly Swin-S and Swin-B, in accurately classifying mosquito species.
Model Precision Recall F1 score Accuracy () Parameters (million) ResNet50 0.937 0.933 0.933 94.50 23.59 ViT Base 0.979 0.978 0.977 97.80 85.81 CvT-13 0.985 0.984 0.984 98.40 19.62 MobileViT 0.989 0.989 0.989 98.90 5.58 Xception 0.991 0.911 0.991 99.20 22.91 Swin-B 1.000 1.000 1.000 99.70 86.75 Swin-S 0.998 0.998 0.998 99.80 48.84
Figure 7 shows the training and validation accuracies and losses vs. epochs for the applied models with closed-set learning. The model demonstrates a consistent rise in training accuracy, with validation accuracy improving, though with some fluctuations. The loss metrics indicate a steady reduction in training and validation loss, highlighting the model’s capability to learn from the data and generalize well. In Figure 7(b), the MobileViT model shows an upward trend in training and validation accuracy, though the validation accuracy increases gradually. Both training and validation loss decrease, but the validation loss plateaus slightly, indicating potential challenges in further improving validation performance. For the ViT Base model, Figure 7(c) illustrates an increasing trend in training and validation accuracy, reflecting effective training. The training loss decreases consistently, while the validation loss reduces, albeit with minor fluctuations, suggesting good generalization with occasional overfitting. According to Figure 7(d), the Xception model indicates a steady rise in training accuracy, with the validation accuracy showing a similar pattern but with slight instability. The training loss decreases significantly, and the validation loss also reduces, though it exhibits minor variations, indicating some challenges in validation performance consistency.
Figure 8 represents the normalized confusion matrix of the applied models for the test samples of the ten mosquito classes. According to Figure 8(a), the Swin Transformer model demonstrates high accuracy, with most species correctly classified. However, there are minor misclassifications, such as Anopheles arabiensis being confused with Anopheles Coluzzii and Anopheles Atroparvus. The MobileViT model attained almost similar performance with Swin in mosquito classification. This lightweight model with significantly reduced parameters achieves high classification accuracy overall, although some misclassifications are observed, such as Anopheles Arabiensis being occasionally misidentified as Anopheles coluzzii. The ViT Transformer model shows exceptional performance, though some misclassifications, such as Anopheles arabiensis being occasionally misidentified as Anopheles coluzzii and Anopheles albimanus.
Figure 9 illustrates the ROC curves for different models applied to closed-set learning, including Swin-B, Xception, ViT, and Swin-S. The curves demonstrate the high performance of the applied models, with most achieving near-perfect AUC scores, indicating their strong capability in distinguishing between mosquito species.
Figure 10 illustrates the classification results of the Swin-B model, displaying the actual and predicted classes for various mosquito species in a closed-set learning scenario. The model accurately identifies each species, with high confidence scores (1.00) for most predictions, demonstrating its effectiveness in distinguishing between different mosquito species.
4.2 Open-set Learning Implementation Results
The primary challenge we encountered during our research was that our models were susceptible to misclassification when presented with images outside the domain of our dataset. Our models, which were trained exclusively on our mosquito dataset, incorrectly classified images of unknown classes (e.g., ants, bees, butterflies) as one of the ten known mosquito classes. This issue is illustrated in Figure 11, where even images of ants, bees, and butterflies were erroneously predicted to belong to one of the mosquito classes. Figure 11(a) shows that an ant was classified as Aedes albopictus with a confidence score of 0.54. In figure 11(b),11(c), two different types of bee were classified as Aedes aegypti and Culex quinquefasciatus with scores of 0.47 and 0.87, respectively. In figure 11(d),11(e), two different types of butterfly were classified as Aedes aegypti with scores of 0.77 and 0.71, respectively. To address this problem, we incorporated the OpenMax method, which allowed us to effectively identify and label outlier images as unknown, thereby improving the robustness of our model in handling images from outside the training dataset.
The primary challenge encountered in our research was the susceptibility of the applied Swin model to misclassifications when presented with images outside the domain of our dataset. Our models, trained exclusively on mosquito images, incorrectly classified unknown insects, such as ants, bees, and butterflies, as one of the known mosquito classes, as illustrated in Figure 11. These misclassifications are expected in a static environment, as the applied models attained incomplete knowledge during training, leading to reduced robustness and flexibility. For instance, an ant was classified as Aedes albopictus with a confidence score of 0.54, while bees and butterflies were erroneously predicted as various mosquito species. To address this issue, we implemented the open-set learning technique with the OpenMax method, which effectively identified and labeled outlier images as unknown. This enhancement improved the robustness of our model, allowing it to better handle images from outside the training dataset and reducing the misclassification rate.
Figure 12 displays the Weibull distribution plots, obtained from (4), for all ten classes identified by the Swin-S model. Each subplot corresponds to a different mosquito species, demonstrating how the model’s predictions follow a Weibull distribution for each class. The Weibull distributions’ consistent shape and scale parameters suggest that the Swin-S model effectively captures the underlying data distribution for each class, indicating robust performance in classifying these species.
Figure 13 showcases the ROC curves for different models (Swin-B, Xception, ViT, and Swin-S) in the context of open-set learning with OpenMax and Weibull distribution. The visual comparison indicates that OpenMax generally enhances the models’ robustness in open-set scenarios by improving their true positive rates while managing false positives. Each class, including the micro-average, achieves an Area Under the Curve (AUC) of 1.00, indicating perfect classification performance. This condition suggests that the model is highly effective at distinguishing between known and unknown classes, demonstrating its robust capability in an open-set recognition scenario.
Model Threshold Accuracy F1 Score Macro Average Weighted Average Precision Recall F1 Score Precision Recall F1 Score Swin-B 0.5 0.54 0.54 0.92 0.54 0.63 0.92 0.54 0.63 0.4 0.65 0.65 0.93 0.65 0.72 0.93 0.65 0.72 0.3 0.73 0.73 0.93 0.73 0.79 0.93 0.73 0.79 0.2 0.81 0.81 0.94 0.81 0.85 0.94 0.81 0.85 0.1 0.88 0.88 0.95 0.88 0.90 0.95 0.88 0.90 Xception 0.5 0.60 0.60 0.93 0.60 0.68 0.93 0.60 0.68 0.4 0.68 0.68 0.93 0.68 0.75 0.93 0.68 0.75 0.3 0.77 0.77 0.93 0.77 0.81 0.93 0.77 0.81 0.2 0.84 0.84 0.94 0.84 0.87 0.94 0.84 0.87 0.1 0.90 0.90 0.94 0.90 0.91 0.94 0.90 0.91 ViT 0.5 0.53 0.53 0.92 0.53 0.62 0.92 0.53 0.62 0.4 0.64 0.64 0.92 0.64 0.71 0.92 0.64 0.71 0.3 0.74 0.74 0.93 0.74 0.89 0.93 0.74 0.89 0.2 0.82 0.82 0.93 0.82 0.85 0.93 0.82 0.85 0.1 0.89 0.89 0.94 0.89 0.91 0.94 0.89 0.91 ResNet50 0.5 0.57 0.57 0.86 0.57 0.64 0.86 0.57 0.64 0.4 0.66 0.66 0.86 0.66 0.71 0.86 0.66 0.71 0.3 0.74 0.74 0.86 0.74 0.77 0.86 0.77 0.74 0.2 0.79 0.79 0.87 0.79 0.81 0.87 0.79 0.81 0.1 0.85 0.85 0.88 0.85 0.86 0.88 0.85 0.86 Swin-S 0.5 0.58 0.58 0.93 0.58 0.66 0.93 0.58 0.66 0.4 0.68 0.68 0.93 0.68 0.75 0.93 0.68 0.75 0.3 0.76 0.76 0.93 0.76 0.81 0.93 0.76 0.81 0.2 0.83 0.83 0.94 0.83 0.86 0.94 0.83 0.86 0.1 0.89 0.89 0.95 0.89 0.90 0.95 0.89 0.90
Table 3 presents the performance metrics of the applied models (Swin-B, Xception, ViT, ResNet50, and Swin-S) across various thresholds, calculated from (6), while implementing the OpenMax technique. Among the models, Xception demonstrates the best performance, achieving the highest accuracy of 90% and an F1 score of 0.90 at the 0.10 threshold along with the highest macro and weighted average F1 score of 0.91, indicating its superior capability in handling open-set recognition tasks compared to the other models.
Figure 14 depicts the relationship between OpenMax threshold values and accuracy for various models, including Swin-B, Swin-S, ResNet50, ViT, and Xception. Among these models, the Xception model exhibits the highest accuracy at a threshold value of 0.10, achieving an accuracy of 90%. As the threshold value increases from 0.10 to 0.50, all models show a general decline in accuracy, with the Xception model consistently performing better at lower thresholds.
Figure 15 illustrates the classification results of the Xception model, which performed the best among the models enhanced with the OpenMax technique on an open-set dataset comprising 11 classes, including both known (10 mosquito species) and unknown categories. Each subfigure depicts an image from the dataset with its corresponding actual and predicted class labels, highlighting the model’s ability to handle open-set recognition by identifying classes not seen during training, such as “Human”, “Butterfly”, “Bee”, and “Fly” as “Unknown”. This situation demonstrates the effectiveness of OpenMax in developing the robustness of the Xception model by enabling it to recognize and manage previously unseen classes, thus reducing the likelihood of misclassification in a real-world scenario.
Study Dataset Number of Classes Best Model Accuracy (%) F1 Score [6] IEEE data port 2 ETC 99.20% 0.99 [7] Open-source, 2400 images 2 VGG16 98.25% 0.95 [8] Custom, 5400 images 6 VGG16 91.17% N/A [11] 9900 images 7 Swin Transformer 96.30% 0.98 [12] Custom, 397 images 2 QSVM 93% N/A [15] N/A 16 Xception 97.040.87% 0.97 [16] 2648 images 3 GoogLeNet 96% N/A Proposed Merged 4 public datasets 5000 images 10 Swin-B Transformer 99.70% 1.0
Table 4 compares the proposed mosquito classification system with related studies. The proposed system, which merges four public datasets totaling 5000 images and ten mosquito classes and uses the Swin-B model, achieves the highest accuracy of 99.7%.
5 Conclusion
In this research, transformer-based deep learning techniques have been applied for mosquito classification. A comprehensive dataset comprising ten individual mosquito species (two, seven and one types of Aedes, Anopheles and Culex, respectively) has been constructed from public repositories utilized in related articles. Standard preprocessing and augmentation techniques have been implemented to balance the dataset. The Swin transformer model attains the best performance for closed-set learning in static settings, i.e., test/inference samples’ classes have been included in the training set. The lightweight MobileVit technique offers almost equal performance with significantly reduced model complexity and parameter number. The robustness of the applied classification models has been enhanced by applying an open-set learning approach. The Xception deep learning model with the optimized OpenMax threshold demonstrates the best performance in open-set learning. A possible extension of this automatic mosquito classification system can be developing a diverse dataset with a wider variety of mosquito species and genera. Additionally, incorporating more sophisticated open-set and en-world learning techniques, e.g., MetaMax, may improve the system’s ability to handle unseen classes, thus increasing its practical applicability in real-world scenarios. In the future, mosquito images and wingbeat sounds with different backgrounds and noises can be merged to develop a multimodal dataset. The proposed classification system can be deployed into smartphones or IoT devices for real-time detection, increasing community participation.
Conflicts of Interest
No conflicts of interest.
Data Availability
The dataset and code of this work can be found at
Advanced-Vision-Transformers-and-Open-Set-Learning-for-Mosquito-Classification